# Virtual Target Trajectory Prediction for Stochastic Targets

### v1

@@ Abstract
Trajectory prediction of other vehicles is crucial for autonomous vehicles, with applications from missile guidance to UAV collision avoidance. Typically, target trajectories are assumed deterministic, but real-world aerial vehicles exhibit stochastic behavior, such as evasive maneuvers or gliders circling in thermals. This paper uses Conditional Normalizing Flows, an unsupervised Machine Learning technique, to learn and predict the stochastic behavior of targets of guided missiles using trajectory data. The trained model predicts the distribution of future target positions based on initial conditions and parameters of the dynamics. Samples from this distribution are clustered using a time series k-means algorithm to generate representative trajectories, termed virtual targets. The method is fast and target-agnostic, requiring only training data in the form of target trajectories. Thus, it serves as a drop-in replacement for deterministic trajectory predictions in guidance laws and path planning. Simulated scenarios demonstrate the approach’s effectiveness for aerial vehicles with random maneuvers, bridging the gap between deterministic predictions and stochastic reality, advancing guidance and control algorithms for autonomous vehicles.
@@ 1 Introduction
Many traditional prediction models for target motion assume deterministic behavior, either implicitly or explicitly. For example, the Proportional Navigation (PN) guidance law [1] assumes that the target maintains a constant velocity, while the Zero-Effort-Miss (ZEM) guidance law [2, 3] allows for arbitrary target dynamics within a deterministic framework.
@@
However, relying on a deterministic model for the target can degrade the performance of the guidance law. This is because the control commands are based on a fixed prediction of the trajectory of the target, which may not align with the actual, more varied maneuvers that the target is executing. In reality, the motion of the target is not confined to a single deterministic model; it is capable of executing a wide range of possible maneuvers. Recognizing this limitation, [4] introduced an approach using an Extended Kalman Filter (EKF) to estimate not only the position and velocity of the target, but also the parameters governing its dynamics.
This concept was further advanced in [5], where a multi-hypothesis guidance approach was proposed. This method employed an Interacting Multiple Model (IMM) filter to estimate both the state of the target and the probability of various dynamics hypotheses being true at any given time. The IMM framework runs multiple EKFs in parallel, each corresponding to a different target dynamics hypothesis. In turn, multiple guidance laws are executed in parallel, with the final control commands derived by combining the outputs of the guidance laws according to the probability estimates of the various hypotheses.
While this multi-hypothesis guidance technique is powerful, it has two significant drawbacks. First, the number of hypotheses that can be considered is limited by the available computational resources, as each additional hypothesis requires one additional parallel EKF and one additional guidance law. Second, despite incorporating multiple hypotheses, the approach still assumes that the target follows one of several deterministic trajectories, which may not fully capture the stochastic nature of the motion of the target.
To address these limitations, this paper introduces a probabilistic approach. Instead of assuming that the target follows one of several fixed trajectories, we model the maneuvers of the target as a continuous probability distribution. This shift towards a probabilistic framework presents new challenges, particularly in the prediction of the future trajectory of the target, which is now described by a dynamic probability distribution rather than a single deterministic path.
In [6], a Monte Carlo method is used to approximate the distribution of future positions of a target, but due to the computational demands of the method, only a few samples can be generated.
@@
In general, the future states of the target cannot be described as distinct trajectories, but rather as a time-varying probability distribution over the state space. Such problems are called multimodal trajectory prediction problems and have mostly been studied in the context of autonomous driving and pedestrians, but not in the context of guided missiles. One challenge for this kind of problem is that the probability distribution over the state space usually cannot be derived analytically, but rather has to be approximated in some way.
Ref. [7] gives a good overview of the different approaches that have been taken to solve this problem. The methods range from adding noise to a deterministic prediction over anchor methods [8], where likely end points of the trajectories are used to generate multiple trajectories, clustering and Gaussian Mixture Model approaches [9], over grid-based methods [10] to generative models Machine Learning (ML) techniques. One example of ML approaches is Generative Adversarial Networks (GANs) [11], where a classifier is trained to distinguish between real data and data generated by a generator network. In a minimax game, the generator network is trained to generate trajectories that are indistinguishable from real trajectories.
Another ML approach is Variational Autoencoders (VAEs) [12], where an encoder network is trained to encode the input trajectory into a latent space, which is then decoded by a decoder network to generate a trajectory.
@@
Lastly, Normalizing Flows (NFs) [13, 14] are a class of generative ML models that can be used to approximate probability distributions by transforming a simple base distribution into the desired, complex distribution and vice versa using a series of invertible transformations. Compared to other ML approaches like GANs or VAEs, NFs have the advantage that they can perform both sampling (of possible future states) and inference (i.e., calculating the probability of a given state). Moreover, they allow for efficient and exact computation of the probability of a given state, giving rise to heatmap-like visualizations of the probability density function (PDF) over the state space.
@@
In [15] Conditional NFs (CNFs) are used to predict the future trajectories of pedestrians in a multimodal way. It employs recurrent neural networks to encode the past trajectory of the pedestrians which is used to calculate the conditional probability distribution over the future trajectory conditioned on the past trajectory. Each predicted trajectory consists of a sequence of positions with equal time intervals between them. In [16] a method is proposed to improve the quality and diversity of trajectories generated by NFs. By adding a diversity objective function, more diverse predictions for the future trajectory of vehicles with discrete time steps can be predicted conditioned on measurements and additional physical attributes. To accommodate the need for fast inference, [17] proposes a method to speed up the inference of NFs by reusing the results of previous computations to predict future trajectories of humans.
@@
While most prediction methods aim to predict the future positions for discrete time steps, only the method presented in our previous work [18] is able to predict the probability distribution of the future positions for arbitrary times by using CNFs. With the CNFs approach, the distribution of the future position of stochastically moving targets can be predicted. The advantage of this approach is that the predictions in the form of a PDF can be interpreted and visualized as a heatmap, which can be used to gain insights into the predicted behavior of the target. However, almost all algorithms (guidance laws, collision avoidance, …) expect deterministic trajectories, which is a simplification that might not hold in reality. To bridge this gap, this paper employs a clustering algorithm to cluster the samples generated by the CNFs into a set of representative trajectories, called virtual targets. Since the samples can be generated for equal time steps with equal sample sizes, most of the problems that arise when clustering trajectory data can be avoided. Thus, a time series k-means clustering algorithm [19] is used to generate trajectories of the virtual targets. These virtual targets can then be used as a drop-in replacement for deterministic trajectory predictions in guidance laws, path planning, or other applications that require deterministic trajectory predictions.
This paper leverages the data-driven approach of Conditional Normalizing Flows to learn the stochastic behavior of an aerial vehicle represented by a time-varying PDF of the future positions of the target. The contributions of this paper are two-fold:
@@
    A prediction framework leveraging CNFs to predict the distribution of the future positions of a stochastically moving target, allowing for interpretable, target-agnostic, and fast predictions.
@@
    A time series k-means clustering algorithm is described to cluster the superimposed samples generated by the CNFs for each real target into a set of representative trajectories, called virtual target trajectories, which can be used as a drop-in replacement for deterministic trajectory predictions for various applications.
@@ 3 Learning Stochastic Target Dynamics
The first step to solve the problem described in Sec. 2 is to learn the stochastic behavior of the target. In order to learn the probability distribution of the target’s future position, NFs are used. NFs are a class of generative ML models that can be used to approximate probability distributions by transforming a simple base distribution into the desired, complex distribution and vice versa. Here, the complex distribution is the distribution of the position of the target at a given time and the base distribution is a Gaussian distribution. In the following, the theory of NFs is explained and their application to the problem at hand is described.
@@ 3.1 Normalizing Flows
NFs consist of a series of invertible transformations, which transform a sample from the base distribution into a sample from the complex distribution and vice versa. Since the transformations are invertible, the model can be used for both sampling and inference. Sampling means that the NFs model can be used to generate samples from the desired distribution by sampling from the base distribution and transforming the samples with the learned invertible transformations. Inference, or density estimation, means that the model can be used to calculate the probability density of a given sample from the complex distribution by transforming it to the base distribution and calculating the probability of the sample in the known base distribution and correcting for the volume change due to the transformation by multiplying with the determinant of the Jacobian of the transformation.
As displayed in Fig. 2, NFs describe a transformation $\mathbf{f}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, which transforms samples $\mathbf{x}$ with dimension $d$ from the complex distribution $p(\mathbf{x})$ into samples $\mathbf{z}$ with the same dimension from the base distribution $p(\mathbf{z})$. In this application, $p(\mathbf{z})$ is a Gaussian distribution with zero mean and unit variance and $p(\mathbf{x})$ denotes the distribution of the position $\mathbf{x}$ of the target.
Equation (2) describes the transformation $\mathbf{f}$ and its inverse $\mathbf{f}^{-1}$ as a composition of $n_{\mathrm{l}}$ invertible transformations $\mathbf{f}_{\Theta_{i}}$, where $\Theta_{i}$ are the $n_{\Theta_{i}}$ parameters of the $i$-th transformation. The transformations $\mathbf{f}_{\Theta_{i}}$ can be any invertible function, provided they are easy to evaluate, invert, and differentiate. Moreover, the determinant of the Jacobian of the transformation $\mathbf{f}_{\Theta_{i}}$ must be easy to compute since it is needed for the calculation of the PDF of the complex distribution. Examples of such transformations include affine transformations (e.g., $\mathbf{f}_{\Theta_{i}}(\mathbf{x})=\mathbf{A}\mathbf{x}+\mathbf{b}$), as utilized in RealNVP [20]. In this paper, $\mathbf{f}_{\Theta_{i}}$ are rational quadratic spline transformations, which are elaborated upon in Sec. 3.3.
The parameters $\Theta_{i}$ are the outputs of neural networks (NNs) $\alpha_{i}$, which are trained to learn the parameters of the transformations.
@@
Equation (3) depicts the input and output dimensions of the NNs $\alpha_{i}$ with $d$ being the dimension of the input data and $n_{\Theta_{i}}$ being the size of the parameters of the transformation $\mathbf{f}_{\Theta_{i}}$. The input for $\alpha_{i}$ is the output from the previous transformation $\mathbf{f}_{\Theta_{i-1}}$ as illustrated in Fig. 3.
@@
Figure 2: Illustration of the transformation of points from a normal distribution to a complex distribution (and vice versa) using Normalizing Flows. Their relative probability is according to the probability density function above. Figure adapted from [18].
Figure 3: Illustration of the Normalizing Flows concept with neural networks $\alpha_{i}$ and transformations $f_{\Theta_{i}}$ with parameters $\Theta_{i}$. Figure adapted from [18].
@@ 3.2 Conditional Normalizing Flows
As displayed in Fig. 4, CNFs describe a transformation $\mathbf{f}$, which transforms samples $\mathbf{x}$ from the complex distribution $p(\mathbf{x}\mid t,\mathbf{\psi})$ into samples $\mathbf{z}$ from the base distribution $p(\mathbf{z})$.
Equation (5) describes the transformation $\mathbf{f}$ and its inverse $\mathbf{f}^{-1}$ as a composition of $n_{\mathrm{l}}$ invertible transformations $\mathbf{f}_{\Theta_{i}}$, where $\Theta_{i}$ are the $n_{\Theta_{i}}$ parameters of the $i$-th transformation. The parameters $\Theta_{i}$ are the outputs of NNs $\beta_{i}$, which are trained to learn the parameters of the transformations conditioned on $t$ and $\mathbf{\psi}$.
@@ 3.3 Application of Conditional Normalizing Flows
Leveraging CNFs allows to transform samples from the complex distribution to the base distribution for a given time $t$ and additional parameters $\mathbf{\psi}$ regarding the target dynamics which serve as the conditioning variables. Thus, the PDF $p(\mathbf{x}\mid t,\mathbf{x}_{0},\mathbf{\psi})$ can be evaluated by transforming a position $\mathbf{x}$ from the complex distribution to the base distribution and calculating the probability of the sample in the base distribution. Moreover, the model can be used to sample from the complex distribution for a given time $t$ and parameters $\mathbf{\psi}$ by sampling from the base distribution and transforming the samples with the learned invertible transformations. The calculation time for the evaluation of the PDF or sampling is constant, since the model learns the distribution of the target’s future position and not the dynamics of the target itself. Not only is the sampling using the CNFs much faster than Monte Carlo simulation (especially for evaluations at the final time), the use of CNFs also allows for an exact computation of the PDF, which is not possible with the Monte Carlo trajectory generation method. To facilitate the training process, the training data is normalized to a range of $[-1,1]$, including the positions $\mathbf{x}$, the time $t$, and the parameters $\mathbf{\psi}$.

### v2

@@ Abstract
Trajectory prediction of aerial vehicles is a key requirement in applications ranging from missile guidance to UAV collision avoidance. While most prediction methods assume deterministic target motion, real-world targets often exhibit stochastic behaviors such as evasive maneuvers or random gliding patterns. This paper introduces a probabilistic framework based on Conditional Normalizing Flows (CNFs) to model and predict such stochastic dynamics directly from trajectory data. The learned model generates probability distributions of future target positions conditioned on initial states and dynamic parameters, enabling efficient sampling and exact density evaluation. To provide deterministic surrogates compatible with existing guidance and planning algorithms, sampled trajectories are clustered using a time series k-means approach, yielding a set of representative “virtual target” trajectories. The method is target-agnostic, computationally efficient, and requires only trajectory data for training, making it suitable as a drop-in replacement for deterministic predictors. Simulated scenarios with maneuvering and ballistic targets demonstrate that the proposed approach bridges the gap between deterministic assumptions and stochastic reality, advancing guidance and control algorithms for autonomous vehicles.
44footnotetext: Substantial parts presented as Paper CEAS-GNC-2024-103 at the CEAS EuroGNC 2024, Bristol, UK, 11-13 June 2024.
@@ 1 Introduction
Traditional prediction models for target motion assume deterministic behavior, either implicitly or explicitly. However, relying on a deterministic model for the target can degrade the performance of the guidance law because the control commands are based on a fixed prediction of the trajectory that may not align with the actual, more varied maneuvers that the target is executing. In reality, the motion of the target is not confined to a single deterministic model; it is capable of executing a wide range of possible maneuvers.
@@
This shift towards a probabilistic framework presents new challenges, particularly in the prediction of the future trajectory of the target, which is now described by a dynamic probability distribution rather than a single deterministic path. The future states of the target cannot be described as distinct trajectories, but rather as a time-varying probability distribution over the state space. Such problems are called multimodal trajectory prediction problems and have mostly been studied in the context of autonomous driving and pedestrians, but not in the context of guided missiles.
One challenge for this kind of problem is that the probability distribution over the state space usually cannot be derived analytically, but rather has to be approximated in some way. Moreover, most algorithms (guidance laws, collision avoidance, etc.) expect deterministic trajectories, which creates a gap between the probabilistic nature of target motion and the requirements of downstream applications.
Many traditional approaches have attempted to address target motion prediction. For example, the Proportional Navigation (PN) guidance law [1] assumes that the target maintains a constant velocity, while the Zero-Effort-Miss (ZEM) guidance law [2, 3] allows for arbitrary target dynamics within a deterministic framework.
Recognizing the limitations of purely deterministic approaches, [4] introduced an approach using an Extended Kalman Filter (EKF) to estimate not only the position and velocity of the target, but also the parameters governing its dynamics. This concept was further advanced in [5], where a multi-hypothesis guidance approach was proposed using an Interacting Multiple Model (IMM) filter to estimate both the state of the target and the probability of various dynamics hypotheses.
While this multi-hypothesis guidance technique is powerful, it has two significant drawbacks. First, the number of hypotheses that can be considered is limited by computational resources. Second, despite incorporating multiple hypotheses, the approach still assumes that the target follows one of several deterministic trajectories, which may not fully capture the stochastic nature of target motion.
In [6], a Monte Carlo method is used to approximate the distribution of future positions of a target, but due to computational demands, only a few samples can be generated. Ref. [7] gives a good overview of different approaches to multimodal trajectory prediction, ranging from adding noise to deterministic predictions, anchor methods [8], clustering and Gaussian Mixture Model approaches [9], grid-based methods [10], to generative Machine Learning (ML) techniques.
ML approaches include Generative Adversarial Networks (GANs) [11], where a classifier is trained to distinguish between real data and data generated by a generator network. In a minimax game, the generator network is trained to generate trajectories that are indistinguishable from real trajectories. Variational Autoencoders (VAEs) [12] allow for the generation of new samples by encoding input data into a latent space and then decoding it back to the original space. Diffusion models [13] are a recent class of generative models that have shown great success in generating high-quality samples. However, diffusion models typically require hundreds of function evaluations for inference, making them impractical for real-time applications like missile guidance. In [14], a diffusion model is compared to other methods for trajectory prediction, performing significantly worse than them.
@@
Normalizing Flows (NFs) [15, 16] are a class of generative ML models that can be used to approximate probability distributions by transforming a simple base distribution into the desired, complex distribution and vice versa using a series of invertible transformations. Compared to other ML approaches like GANs or VAEs, NFs have the advantage that they can perform both sampling (of possible future states) and inference (i.e., calculating the probability of a given state). Moreover, they allow for efficient and exact computation of the probability of a given state, giving rise to heatmap-like visualizations of the probability density function (PDF) over the state space.
@@
In [17] Conditional NFs (CNFs) are used to predict the future trajectories of pedestrians in a multimodal way. It employs recurrent neural networks to encode the past trajectory of the pedestrians which is used to calculate the conditional probability distribution over the future trajectory conditioned on the past trajectory. Each predicted trajectory consists of a sequence of positions with equal time intervals between them.
In [18] a method is proposed to improve the quality and diversity of trajectories generated by NFs. By adding a diversity objective function, more diverse predictions for the future trajectory of vehicles with discrete time steps can be predicted conditioned on measurements and additional physical attributes. To accommodate the need for fast inference, [19] proposes a method to speed up the inference of NFs by reusing the results of previous computations to predict future trajectories of humans.
@@
Only the method in our previous work [20] predicts probability distributions of future positions for arbitrary times using CNFs. With the CNFs approach, the distribution of the future position of stochastically moving targets can be predicted. The advantage of this approach is that the predictions in the form of a PDF can be interpreted and visualized as a heatmap, which can be used to gain insights into the predicted behavior of the target. However, almost all algorithms (guidance laws, collision avoidance, …) expect deterministic trajectories, which is a simplification that might not hold in reality. To bridge this gap, this paper employs a clustering algorithm to cluster the samples generated by the CNFs into a set of representative trajectories, called virtual targets. Since the samples can be generated for equal time steps with equal sample sizes, most of the problems that arise when clustering trajectory data can be avoided. Thus, a time series k-means clustering algorithm [21] is used to generate trajectories of the virtual targets. These virtual targets can then be used as a drop-in replacement for deterministic trajectory predictions in guidance laws, path planning, or other applications that require deterministic trajectory predictions. The contributions of this paper are two-fold:
@@
    A prediction framework leveraging CNFs to predict the distribution of future positions of stochastically moving targets, allowing for interpretable, target-agnostic, and fast predictions.
@@
    A time series k-means clustering algorithm to cluster samples generated by CNFs into representative trajectories, called virtual target trajectories, which can be used as a drop-in replacement for deterministic trajectory predictions for various applications.
@@ 3 Learning Stochastic Target Dynamics
The first step to solve the problem described in Sec. 2 is to learn the stochastic behavior of the target. In order to learn the probability distribution of the target’s future position, NFs are used. NFs are a class of generative ML models that can be used to approximate probability distributions by transforming a simple base distribution into the desired, complex distribution and vice versa. Here, the complex distribution is the distribution of the position of the target at a given time and the base distribution is a Gaussian distribution. In the following, we first provide background on the established theory of NFs and CNFs, before describing their novel application to the problem at hand.
@@ 3.1 Normalizing Flows
NFs are an established class of generative ML models introduced by [15] that consist of a series of invertible transformations, which transform a sample from the base distribution into a sample from the complex distribution and vice versa. They consist of a series of invertible transformations, which transform a sample from the base distribution into a sample from the complex distribution and vice versa. Since the transformations are invertible, the model can be used for both sampling and inference. Sampling means that the NFs model can be used to generate samples from the desired distribution by sampling from the base distribution and transforming the samples with the learned invertible transformations. Inference, or density estimation, means that the model can be used to calculate the probability density of a given sample from the complex distribution by transforming it to the base distribution and calculating the probability of the sample in the known base distribution and correcting for the volume change due to the transformation by multiplying with the determinant of the Jacobian of the transformation.
Figure 2 (adapted from [20]) illustrates the transformation of samples from a normal distribution to a complex distribution (and vice versa) using NFs. The samples are depicted by the shapes (squares, circles, …) and their relative probability is indicated by the probability density function above. In general, NFs describe a bijection $\mathbf{f}$, which transforms samples $\mathbf{x}\in\mathbb{R}^{d}$ from the complex distribution $p(\mathbf{x})$ into samples $\mathbf{z}\in\mathbb{R}^{d}$ from the base distribution $p(\mathbf{z})$. In this application, $p(\mathbf{z})$ is a Gaussian distribution with zero mean and unit variance and $p(\mathbf{x})$ denotes the distribution of the position $\mathbf{x}$ of the target.
Equation (2) describes the transformation $\mathbf{f}$ and its inverse $\mathbf{f}^{-1}$ as a composition of $n_{\mathrm{l}}$ invertible transformations $\mathbf{f}_{\Theta_{i}}$, where $\Theta_{i}$ are the $n_{\Theta_{i}}$ parameters of the $i$-th transformation. The transformations $\mathbf{f}_{\Theta_{i}}$ can be any invertible function, provided they are easy to evaluate, invert, and differentiate. Moreover, the determinant of the Jacobian of the transformation $\mathbf{f}_{\Theta_{i}}$ must be easy to compute since it is needed for the calculation of the PDF of the complex distribution. Examples of such transformations include affine transformations (e.g., $\mathbf{f}_{\Theta_{i}}(\mathbf{x})=\mathbf{A}\mathbf{x}+\mathbf{b}$), as utilized in RealNVP [22]. In this paper, $\mathbf{f}_{\Theta_{i}}$ are rational quadratic spline transformations, which are elaborated upon in Sec. 3.3. The parameters $\Theta_{i}$ are the outputs of neural networks (NNs) $\alpha_{i}$, which are trained to learn the parameters of the transformations.
@@
Equation (3) depicts the input and output dimensions of the NNs $\alpha_{i}$ with $d$ being the dimension of the input data and $n_{\Theta_{i}}$ being the size of the parameters of the transformation $\mathbf{f}_{\Theta_{i}}$. The input for $\alpha_{i}$ is the output from the previous transformation $\mathbf{f}_{\Theta_{i-1}}$ as illustrated in Fig. 3 (adapted from [20]).
@@
Figure 2: Illustration of the transformation of samples from a normal distribution to a complex distribution (and vice versa) using Normalizing Flows. Figure adapted from Schneider, M., Loureiro, R., Cunis, T., and Fichter, W., “Trajectory Prediction for Missile Targets: A Probabilistic Approach Using Machine Learning,” in CEAS EuroGNC Conference, 2024. Licensed under a Creative Commons Attribution 4.0 International License ($C$-BY 4.0).
Figure 3: Illustration of the Normalizing Flows concept with neural networks $\alpha_{i}$ and transformations $f_{\Theta_{i}}$ with parameters $\Theta_{i}$. Figure adapted from Schneider, M., Loureiro, R., Cunis, T., and Fichter, W., “Trajectory Prediction for Missile Targets: A Probabilistic Approach Using Machine Learning,” in CEAS EuroGNC Conference, 2024. Licensed under a Creative Commons Attribution 4.0 International License ($C$-BY 4.0).
@@ 3.2 Conditional Normalizing Flows
As displayed in Fig. 4, CNFs describe a transformation $\mathbf{f}$, which transforms samples $\mathbf{x}$ from the complex distribution $p(\mathbf{x}\mid t,\mathbf{\psi})$ into samples $\mathbf{z}$ from the base distribution $p(\mathbf{z})$. Equation (5) describes the transformation $\mathbf{f}$ and its inverse $\mathbf{f}^{-1}$ as a composition of $n_{\mathrm{l}}$ invertible transformations $\mathbf{f}_{\Theta_{i}}$, where $\Theta_{i}$ are the $n_{\Theta_{i}}$ parameters of the $i$-th transformation. The parameters $\Theta_{i}$ are the outputs of NNs $\beta_{i}$, which are trained to learn the parameters of the transformations conditioned on $t$ and $\mathbf{\psi}$.
@@ 3.3 Application of Conditional Normalizing Flows
Leveraging CNFs allows to transform samples from the complex distribution to the base distribution for a given time $t$ and additional parameters $\mathbf{\psi}$ regarding the target dynamics which serve as the conditioning variables. Thus, the PDF $p(\mathbf{x}\mid t,\mathbf{x}_{0},\mathbf{\psi})$ can be evaluated by transforming a position $\mathbf{x}$ from the complex distribution to the base distribution and calculating the probability of the sample in the base distribution. Moreover, the model can be used to sample from the complex distribution for a given time $t$ and parameters $\mathbf{\psi}$ by sampling from the base distribution and transforming the samples with the learned invertible transformations. The calculation time for the evaluation of the PDF or sampling is constant, since the model learns the distribution of the target’s future position and not the dynamics of the target itself. Not only is the sampling using the CNFs much faster than Monte Carlo simulation (especially for evaluations at the final time), the use of CNFs also allows for an exact computation of the PDF, which is not possible with the Monte Carlo trajectory generation method. To facilitate the training process, the training data is normalized to a range of [1,1][-1,1], including the positions $\mathbf{x}$, the time $t$, and the parameters $\mathbf{\psi}$.